OpenStudy (kittykitty04):

Solving for x: x=sqrt((8.314j/mol*k)(298K)/(0.03995kg/mol)) I have got an answer for this question but I want to know if I'm doing this correctly.

OpenStudy (kittykitty04):

\[\sqrt{\frac{ (8.314\frac{ J }{ mol*k })*(298K) }{ (0.03995kg/mol) }}\]

OpenStudy (kittykitty04):

that is the equation and we are trying to find 'x'

OpenStudy (kittykitty04):

My answer came out to 249.0J/kg does that look correct? Much appreciated for the help!

OpenStudy (jfraser):

your number is fine, but when you look at the units carefully, you have \[\sqrt{\frac{(\frac{J}{mol*K})*(K)}{\frac{kg}{mol}}}\] a joule is a \[\frac{kg*m^2}{s^2}\]and when you take all the units, then cancel the square root, you're left with \[\frac{m}{s}\]which is a speed. That equation is used to calculate the average velocity of a sample of gas, which in this case is around 249m/s

OpenStudy (kittykitty04):

Thank you so much!!